OpenStudy (ksaimouli):

3sin2θ+7sinθ+2=0 slove for theta

OpenStudy (anonymous):

I think you have to find theta such that sin theta=-1/2 and sin2theta=1/2

OpenStudy (anonymous):

Is that: \[\sin(2 \theta) or \sin^2(\theta)\]